Show that ((7+3sqrt5)/(sqrt2+sqrt7+3sqrt5)+((7-3sqrt5)/(sqrt2+sqrt7-3sqrt5))=2sqrt2


1
Expert's answer
2021-03-18T14:27:58-0400

Perhaps, condition is incorrect:

"\dfrac{(7+3\sqrt5)}{(\sqrt2+\sqrt7+3\sqrt5)}+\dfrac{(7-3\sqrt5)}{(\sqrt2+\sqrt7-3\sqrt5)}=2\sqrt2"


"\dfrac{(7+3\sqrt5)}{(\sqrt2+\sqrt7+3\sqrt5)}+\dfrac{(7-3\sqrt5)}{(\sqrt2+\sqrt7-3\sqrt5)}\approx" "1.162845244"


"2\sqrt2\approx" "2.828427124"